# Cross-modal Retrieval Models for Stripped Binary Analysis

### v1

@@ Abstract
LLM-agent based binary code analysis has demonstrated significant potential across a wide range of software security scenarios, including vulnerability detection, malware analysis, etc. In agent workflow, however, retrieving the positive from thousands of stripped binary functions based on user query remains under-studied and challenging, as the absence of symbolic information distinguishes it from source code retrieval. In this paper, we introduce, BinSeek, the first two-stage cross-modal retrieval framework for stripped binary code analysis. It consists of two models: BinSeek-Embedding is trained on large-scale dataset to learn the semantic relevance of the binary code and the natural language description, furthermore, BinSeek-Reranker learns to carefully judge the relevance of the candidate code to the description with context augmentation. To this end, we built an LLM-based data synthesis pipeline to automate training construction, also deriving a domain benchmark for future research. Our evaluation results show that BinSeek achieved the state-of-the-art performance, surpassing the the same scale models by 31.42% in Rec@3 and 27.17% in MRR@3, as well as leading the advanced general-purpose models that have 16 times larger parameters.
@@
Guoqiang Chen1,2, Lingyun Ying2, Ziyang Song2, Daguang Liu2, Qiang Wang2, Zhiqi Wang2, Li Hu1 Shaoyin Cheng1, Weiming Zhang1, Nenghai Yu1 1University of Science and Technology of China, 2QI-ANXIN Technology Research Institute
@@ 1 Introduction
Binary code analysis serves as a cornerstone of software security, supporting critical tasks such as vulnerability detection, malware analysis, program audit, and so on. Traditional manual analysis is known to be labor-intensive and requires deep expertise, especially for binaries including thousands of functions. Fortunately, LLM agent system is very promising as the next step in automating binary code analysis, given that LLMs have already demonstrated impressive capabilities in understanding the binary code Shang et al. (2025b, 2024). The agent system takes user queries as input and uses LLMs to plan and execute a series of actions to finally generate the answer, reports or proof-of-conception scripts typically. However, a critical bottleneck impedes further deployment: there is a lack of effective methods to retrieve information related to the user query from a large-scale binary codebase.
@@
Retrieving stripped binary code based on natural language (NL) queries, however, presents unique and formidable challenges compared to source code retrieval. Unlike high-level programming languages (PL), stripped binaries lack explicit semantic indicators such as variable names, function names, and type information. These comprehensive symbols are generally stripped before releasing the binary programs for different reasons, for example, reducing file size, hiding functionality, etc. This absence significantly complicates the understanding of binary code for both humans and models, downgrading the performance of general retrieval models in this scenario Jin et al. (2023).
@@
The problem of semantic-based code search has been extensively studied in the domain of source code. Existing approaches typically learn a joint embedding space to align natural language and code, with pretrained models such as CodeBERT Feng et al. (2020), GraphCodeBERT Guo et al. (2021), and UniXcoder Guo et al. (2022) achieving remarkable success. Despite these advancements, directly adapting source code retrieval models to the binary domain is ineffective due to the fundamental modality differences discussed above. Moreover, unlike the abundance of open-source repositories (e.g., GitHub) that fuel source code representation learning, there is a distinct scarcity of high-quality, aligned datasets, which further hinders the development of robust binary code retrieval models.
@@
Since retrieving binary code with NL queries is rarely studied, in this work, we aim to explore an effective retrieval method for stripped binary analysis. In particular, we introduce BinSeek, the first cross-modal retrieval framework specifically tailored for this issue. A two-stage retrieval strategy is employed to balance accuracy and efficiency: for the first stage, we build, BinSeek-Embedding, a embedding model to retrieve the candidates from a codebase that is decompiled from binary files, and for the second stage, we further devise, BinSeek-Reranker, to reorder the candidates augmented with calling context information for more precise results. To train our expert models, we employ LLMs to automatically synthesize high-quality semantic labels in NL for binary functions. In this way, we also deliver the first domain benchmark for the binary code retrieval task, which is expected to facilitate future research in this domain. Our BinSeek achieves state-of-the-art (SOTA) performance with Rec@3 of 84.5% and MRR@3 of 80.25%, indicating its effectiveness in finding semantic-related binary functions. Our main contributions are summarized as follows:
@@
  • We introduce BinSeek, the first two-stage cross-modal retrieval framework for stripped binary code analysis, where BinSeek-Embedding is used to retrieve candidates from a codebase, and BinSeek-Reranker further refine candidates with calling context information.
@@
  • We propose an LLM-driven data synthesis pipeline to automate the construction of high-quality training data, deriving the first domain benchmark for binary code retrieval.
@@
We make our model and evaluation benchmark publicly available at: https://github.com/XingTuLab/BinSeek.
@@ 2 Background and Challenges
uint32_t v[2], uint32_t const key[4]){
@@
2.1 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) Lewis et al. (2020) searches documents relevant to user queries from external databases, taking them as reference for LLM-based generation, thereby fundamentally expanding the knowledge stored within the model parameters. Since it is highly effective in eliminating model hallucinations and improving generation quality, many efforts have been made to enhance the RAG system, such as GraphRAG Edge et al. (2025), RAGFusion Momii et al. (2024), and KAG Liang et al. (2025).
In software engineering, the retrieval-augmented code generation (RACG) is also an important issue, involving cross-modal retrieval between the PL and NL text. CodeRAG-Bench Wang et al. (2025b) investigated popular embedding models for searching code snippets across the whole repository, and their experiments demonstrated that LLMs benefit from high-quality retrieved results, especially for complex tasks in SWE-Bench Jimenez et al. (2024) (improving pass@1 from 2.3% to 21.7%). Beyond academic research, practical systems such as Cursor Cursor (2025) have developed the codebase indexing to recall highly context-relative code from the codebase. Such retrieval schemes require high-quality code retrieval models to ensure effective performance.
@@ 2.2 Source Code Retrieval Model
Code retrieval models aim to learn a shared embedding space where both NL descriptions and PL code are mapped to dense vector representations, enabling similarity-based retrieval of relevant code snippets. CodeBERT Feng et al. (2020) presented the first bimodal pre-trained model for PL and NL, GraphCodeBERT Guo et al. (2021) further introduced data flow information to enhance code representation learning. UniXcoder Guo et al. (2022) leverages cross-modal contents like abstract syntax tree and code comment to enhance code representation, while CodeXEmbed Liu et al. (2024) designed a family of code embedding models tailored to diverse PLs and retrieval tasks. Furthermore, many comprehensive benchmarks and datasets like CodeSearchNet Husain et al. (2019), CodeXGLUE Lu et al. (2021), and CoIR Li et al. (2024) have facilitated the evaluation of these models, demonstrating their strong capabilities in searching from a large-scale corpus. However, similar research in the binary code domain remains largely unexplored, which motivates our current investigation into cross-modal retrieval models for binary analysis.
2.3 Stripped Binary Code
Binary code is a low-level representation of a program, which is generally compiled from the source code written in PLs like C/C++, and it consists of a sequence of machine instructions encoded in binary format. Due to various reasons (e.g., reducing file size, hiding functionality), most binary files are stripped of symbol information before release, resulting in stripped binary code that lacks function names, variable names, data types, and other high-level semantic information. Although modern decompilers can handle binaries and generate C-like decompiled pseudocode, they must use placeholders to represent missing symbolic information, for example, the function sub_100000A68 and the variable a2/v3 shown in Figure 1.
Recently, many studies have explored automated analysis of binary pseudocode using LLMs, such as LLM4Decompile Tan et al. (2024), ReCopilot Chen et al. (2025), and Bin2Wrong Yang and Nagy (2025), yielding promising experimental results. In practice, however, a binary program usually contains tens of thousands of functions Marcelli et al. (2022); Sun et al. (2023), highlighting the difficulty of searching for specific code within it. It is particularly difficult given the absence of high-level semantic symbols in binary code. Specifically, we summarize the critical challenges as follows:
@@ 2.3 Stripped Binary Code
  • Irreversible Semantics Loss: Given the information loss during compilation and stripping, it is typically impossible to infer the source-level functional semantics of a function solely from its binary code. For instance, the function name in the source code in Figure 1 indicates its operational context for cryptographic encoding, which cannot be deduced from the binary function alone. Therefore, effective binary code retrieval models must learn to infer semantics from the context.
@@ 3 Method
Figure 2 shows the building overview of our BinSeek, which consists of BinSeek-Embedding and BinSeek-Reranker models. BinSeek-Embedding transfer user queries in NL and binary functions in pseudocode into embeddings, assessing the semantic relevance of cross-modal inputs through the distance between these embeddings. Each vector requires only one single model computation, making it suitable for fast retrieval within large corpora. However, considering the challenges mentioned in §2.3, the BinSeek-Embedding model struggles to accurately capture the complete source code-level semantics through a binary function alone, leading to suboptimal retrieval results. To address this, we further introduce the BinSeek-Reranker model as a secondary reranking step. This model takes the binary code with context functions and predicts its relevance to the NL query. By utilizing supplementary information from the calling context, the BinSeek-Reranker model can infer the semantics of the target function more comprehensively, thereby optimizing the retrieval results.
@@ 3.1 Data Synthesis
Training cross-modal retrieval models requires a large amount of data pairs that consist of pseudocode functions and corresponding semantic descriptions in NL. Unfortunately, there is currently no publicly available dataset specifically designed for binary code retrieval. Although hiring reverse engineering experts to label pseudocode is the optimal approach for best data quality, high labor costs make it impractical for large-scale execution. To overcome this, we propose a data synthesis pipeline to automatically generate high-quality training data.
@@
As shown in Figure 3, we first collect a large number of open-source projects written in C/C++ from GitHub. For each project, we compile the source code into stripped binary files using different configurations, while saving the debugging information separately (more details in §A.1). Next, we utilize a modern decompiler (i.e., IDA Pro Hex-RaysSA (2025)) to decompile binary files into pseudocode functions. Since stripped binaries do not have symbols, the debugging information serves as a bridge linking the source code to the binary code in the function level. Finally, we employ an advanced LLM (i.e., DeepSeek DeepSeek-AI et al. (2025)) to generate semantic descriptions for each source code function. More specifically, we provide the LLM with the content of the source code file containing the function, along with the file path, project name, and version number as references (detailed in §A.2). NL descriptions, source code, and binary code are essentially representations of the same semantics across different modalities. We therefore trained our model to learn the mappings between NL descriptions and pseudocode functions. Ultimately, we constructed 106,711,414 raw data pairs in total.
Quality Filters. To ensure the quality of the synthesized data, we design a filter module to eliminate low-quality data pairs. We have designed the following filters:
@@
  • Binary Code Filter removes short binary functions similarly, as well as thunk functions and virtual functions, which typically is the wrapper for jumping or calling other functions.
@@
  • Repetitive Semantics Filter drops both source and binary functions with high similarity to reduce code redundancy. We employ MinHash Broder (1997), a locality-sensitive hashing algorithm, to perform function-level deduplication, and use the empirical threshold of 0.95 Shang et al. (2025a).
Data Sampler. Training retrieval models typically require both positive and negative data pairs. The positive pair naturally consist of the NL description and pseudocode function from the same source code. To construct effective negative samples, we employ a hybrid sampling strategy combining random sampling and hard negative mining. Specifically, for each source code function, we take the NL description as a query and randomly select the binary functions as negative candidates that are compiled from the other source code among the whole corpus. Furthermore, we utilize a general embedding model (Qwen3-Embedding-8B Zhang et al. (2025)), which is top-ranked in the MTEB leaderboard Enevoldsen et al. (2025), to estimate the semantic similarity between the query and the NL descriptions of the negative candidates. We select the binary function as hard negative whose description has a low similarity score ($\leq 0.95$) with the query. In this way, we constructed 45,703,009 high-quality positive and negative data pairs in total, enabling the model to learn the mappings across different semantic representations.
@@ 3.2 Model Architecture
BinSeek-Embedding. The embedding model is built from scratch with reference to the existing language model Qwen3 Yang et al. (2025). Specifically, we initialize our model backbone with an 8 transfomer layer, which incorporates RMSNorm Jiang et al. (2023), SwiGLU activation Dauphin et al. (2017), and grouped query attention (GQA) Ainslie et al. (2023) techniques for improved training stability and performance. We employ Qwen3’s tokenizer that implements byte-level byte-pair encoding Wang et al. (2020); Sennrich et al. (2016) with a vocabulary size of 151,669. Consequently, the weights of the token embedding layer are directly loaded from the well-trained Qwen3 model and frozen to reduce training overhead. Following the embedding layer, we use a rotary position embedding (RoPE) Su et al. (2024) layer to provide position information to transformer blocks, meanwhile enabling input length scaling at inference time. Finally, the last token pooling is employed to the backbone output to generate the final embedding.
@@
To mitigate this, we here propose a heuristic function importance metric to select the top-$k$ important callees to provide context. Existing studies have demonstrated that the presence of meaningful function names and strings often provides significant semantic cues about its functionality Chen et al. (2023); Shang et al. (2025b); Jin et al. (2023). Therefore, we try to measure the meaningful symbol density of a pseudocode function $f_{t}$ through three aspects: (1) whether the function has an unstripped function name, (2) the density of strings among the pseudocode, and (3) the meaningful callee names in the function. Formally, we define the informative score $\mathcal{S}(f_{t})$ to indicate function importance as follows:

### v2

@@ Abstract
Retrieving binary code via natural language queries is a pivotal capability for downstream tasks in the software security domain, such as vulnerability detection and malware analysis. However, it is challenging to identify binary functions semantically relevant to the user query from thousands of candidates, as the absence of symbolic information distinguishes this task from source code retrieval. In this paper, we introduce, BinSeek, a two-stage cross-modal retrieval framework for stripped binary code analysis. It consists of two models: BinSeek-Embedding is trained on large-scale dataset to learn the semantic relevance of the binary code and the natural language description, furthermore, BinSeek-Reranker learns to carefully judge the relevance of the candidate code to the description with context augmentation. To this end, we built an LLM-based data synthesis pipeline to automate training construction, also deriving a domain benchmark for future research. Our evaluation results show that BinSeek achieved the state-of-the-art performance, surpassing the the same scale models by 31.42% in Rec@3 and 27.17% in MRR@3, as well as leading the advanced general-purpose models that have 16 times larger parameters.
@@
Guoqiang Chen1,2, Lingyun Ying2, Ziyang Song2, Daguang Liu2, Qiang Wang2, Zhiqi Wang2, Li Hu1, Shaoyin Cheng1, Weiming Zhang1, Nenghai Yu1 1University of Science and Technology of China, 2QI-ANXIN Technology Research Institute
@@ 1 Introduction
Binary code analysis serves as a cornerstone of software security, supporting critical tasks such as vulnerability detection, malware analysis, program audit. Traditional manual analysis is known to be labor-intensive and requires deep expertise, especially for binaries including thousands of functions. To alleviate these scalability challenges, recent studies BinMetric; shang2024bcodeunderstanding have begun exploring the application of Large Language Models (LLMs) to various binary code-related analysis tasks. Nevertheless, existing work has largely overlooked the problem of retrieving binary code through natural language (NL) queries, leaving a gap in effectively bridging users’ high-level semantic intentions with large-scale binary codebases.
@@
Retrieving stripped binary code with NL queries, however, presents unique and substantial challenges compared to source code retrieval. Unlike high-level programming language (PL), stripped binaries lack explicit semantic indicators such as variable names, function names, and type information. These comprehensive symbols are generally stripped before releasing the binary programs for different reasons (e.g., reducing file size, hiding functionality). This absence significantly complicates the understanding of binary code for both humans and models BinSum-jin2023, downgrading the performance of general-purpose retrieval models in this scenario.
@@
Semantic-based code search has been extensively explored in the context of source code. Existing approaches typically learn a joint embedding space to align NL and code, with pretrained models such as CodeBERT feng2020CodeBERT, GraphCodeBERT guo2021GraphCodeBERT, and UniXcoder guo2022unixcoder achieving remarkable success. Despite these advancements, directly adapting source code retrieval models to the binary domain is ineffective due to the fundamental modality differences discussed above. Moreover, unlike the abundance of open-source repositories (e.g., GitHub) that fuel source code representation learning, a high-quality and well-labeled dataset for binaries remain scarce, constraining the development of robust binary code retrieval models.
@@
Since retrieving binary code with NL queries is rarely studied, this work aims to develop an effective solution for retrieving functions from stripped binaries. In particular, we introduce BinSeek, a two-stage cross-modal retrieval framework specifically tailored for this issue: for the first stage, we build, BinSeek-Embedding, a embedding model to retrieve the candidates from a codebase that is decompiled from binaries, and for the second stage, we further devise, BinSeek-Reranker, to reorder the candidates augmented with calling context information for more precise results. To train our expert models, we employ LLMs to automatically synthesize high-quality semantic labels in NL for binary functions. In this way, we also deliver the first domain benchmark for the binary code retrieval task, which is expected to facilitate future research in this domain. Our BinSeek achieves state-of-the-art (SOTA) performance with Rec@3 of 84.5% and MRR@3 of 80.25%, indicating its effectiveness in finding semantic-related binary functions. Our main contributions are summarized as follows:
@@
  • We introduce BinSeek, a two-stage cross-modal retrieval framework for stripped binary code analysis, where BinSeek-Embedding is used to retrieve candidates from a codebase, and BinSeek-Reranker further refine candidates with calling context information.
@@
  • We propose an LLM-driven data synthesis pipeline to automate the construction of high-quality training data, also deriving the first domain benchmark for binary code retrieval.
@@
We make our model and evaluation benchmark publicly available at here.
@@ 2 Background and Challenges
uint32_t ori[2], uint32_t const key[4]){
@@
2.1 Problem Definition
In this work, we investigate the problem of retrieving binary functions using NL queries. Let $\mathcal{P}$ denote the set of pseudocode decompiled from stripped binary files, where each $p_{i}\in\mathcal{P}$ corresponds to a binary function in the codebase. Given an NL query $q$ describing a desired behavior or functionality, the objective is to search the most semantically relevant pseudocode functions:
$\mathcal{R}(q,\mathcal{P})=\{p_{1},p_{2},\ldots,p_{k}\}$ (1)
This formulation reflects practical scenarios, where human or LLM-based analysts aim to effectively and quickly locate relevant functionality within stripped binaries using high-level descriptions and insights, for tasks such as vulnerability investigation, malware analysis, and software auditing. We focus on decompiled binary code rather than disassembly code, as previous studies have highlighted it is more human-readable and LLM-friendly representation NER-2023pst; BinSum-jin2023.
This task is distinct from binary code similarity detection (BCSD). BCSD methods (e.g., jTrans Wang2022jTrans and CLAP Wang2024CLAP) focuses on measuring semantic similarity between two binary code snippets, typically to determine whether they implement identical or related functionality. In contrast, our task is inherently cross-modal: the input is NL, while the target consists of pseudocode functions. The core challenge lies in bridging the semantic gap between high-level linguistic intent and the semantics implicitly preserved in stripped binary code.
@@ 2.2 Source Code Retrieval Model
Code retrieval models aim to learn a shared embedding space where both NL descriptions and PL code are mapped to dense vector representations, enabling similarity-based retrieval of relevant code snippets. CodeBERT feng2020CodeBERT presented the first bimodal pre-trained model for PL and NL, GraphCodeBERT guo2021GraphCodeBERT further introduced data flow information to enhance code representation learning. UniXcoder guo2022unixcoder leverages cross-modal contents like abstract syntax tree and code comment for more improvement, while CodeXEmbed liu2024codexembed designed a family of code embedding models tailored to diverse PLs and retrieval tasks. Furthermore, many comprehensive benchmarks and datasets like CodeSearchNet husain2019codesearchnet, CodeXGLUE Lu2021CodeXGLUE, and CoIR CoIR have facilitated the evaluation of these models, demonstrating their strong capabilities in searching from a large-scale corpus. However, similar research in the binary code domain remains limited, which motivates our current investigation into cross-modal retrieval models for binary analysis.
2.3 Stripped Binary Code Retrieval
Binary code is a low-level representation of a program, typically compiled from the source code written in PLs like C/C++, and it consists of a sequence of machine instructions encoded in binary format. Due to various reasons (e.g., reducing file size, hiding functionality), most binary files are stripped of symbol information before release, resulting in the lacks of function names, variable names, data types, and other high-level semantic information. Although modern decompilers can handle binaries and generate C-like decompiled pseudocode, they use placeholders to represent missing symbolic information, like the function sub_100000A68 and the variable a2/v3 shown in Figure 1.
Recently, many studies have explored automated analysis of binary pseudocode using LLMs, such as LLM4Decompile tan2024llm4decompile, ReCopilot chen2025recopilot, and Bin2Wrong usenix2025Bin2Wrong, yielding promising experimental results. In practice, however, a binary program usually contains tens of thousands of functions Cisco2022HowMachine; Sun2023GraphMoCo, highlighting the difficulty of searching for specific code within it. To our knowledge, the only related work, BinQuery binquery, conducted investigation on retrieving disassembled code with NL queries (distinct from our problem definition). They demonstrated the potential of learning-based methods, though the performance remain unsatisfactory. It is particularly difficult given the absence of high-level semantic symbols in binary code. Specifically, we summarize the critical challenges as follows:
@@ 2.3 Stripped Binary Code Retrieval
  • Irreversible Semantics Loss: Given the information loss during compilation and stripping, it is typically hard to infer the source-level semantics of a function solely from its binary code. For instance, the function name in the source code in Figure 1 indicates its operational context for cryptographic encoding, which cannot be deduced from the binary function alone. Therefore, effective binary code retrieval models must learn to infer semantics from the context.
@@ 3 Method
Figure 2 illustrates the overall pipeline of BinSeek, which comprises the BinSeek-Embedding and BinSeek-Reranker models. BinSeek-Embedding maps user queries in NL and pseudocode functions into embeddings, measuring cross-modal semantic relevance based on the embedding distances. Each vector requires only one single model computation, making BinSeek-Embedding suitable for fast retrieval over large codebases. However, as discussed in §2.3, BinSeek-Embedding alone may struggle to capture complete semantics from a pseudocode alone, resulting in suboptimal retrieval performance. To mitigate this, we introduce BinSeek-Reranker as a secondary reranking step. BinSeek-Reranker leverages the supplementary information selected from the calling context for a more comprehensive inference of function semantics, thereby optimizing the retrieval results.
@@ 3.1 Data Synthesis
Training cross-modal retrieval models requires a large amount of data pairs that consist of decompiled pseudocode and corresponding semantic descriptions in NL. Unfortunately, there is currently no publicly available dataset specifically designed for binary code retrieval. Although hiring reverse engineering experts to label pseudocode is the optimal approach for the best quality, high labor costs make it impractical for large-scale execution. To overcome it, we propose a data synthesis pipeline to automatically generate high-quality training data.
@@
As shown in Figure 3, we first collect a large number of open-source projects written in C/C++ from GitHub. For each project, we compile the source code into stripped binary files using different configurations, while saving the debugging information separately (more details in §A.1). Next, we utilize a modern decompiler (i.e., IDA Pro IDA) to decompile binary files into pseudocode functions. Since stripped binaries do not have symbols, the debugging information serves as a bridge linking the source code to the binary code in the function level. Finally, we employ an advanced LLM (i.e., DeepSeek 2025deepseekv3) to generate semantic descriptions for each source code function. More specifically, we provide the LLM with the content of the source code file containing the function, along with the file path, project name, and version number as references (detailed in §A.2). NL descriptions, source code, and binary code are essentially representations of the same semantics across different modalities. We therefore synthesize the training data pairs with generated NL descriptions and binary code.
Quality Filters. To ensure the quality of the synthesized data, we eliminate low-quality data pairs with the following filters:
@@
  • Binary Code Filter also removes short binary functions similarly, as well as thunk functions and virtual functions, which typically is the wrapper for calling other functions.
@@
  • Repetitive Semantics Filter drops both source and binary functions with high similarity to reduce code redundancy. We employ MinHash minhash, a locality-sensitive hashing algorithm, to perform function-level deduplication, and use the empirical threshold of 0.95 TOSEM2025FoC.
Ultimately, we constructed 106,711,414 raw data pairs in total. A human evaluation shows that 97.6% of the generated descriptions are rated as correct or at least partially correct, indicating the reliability of the synthesized labels (detailed in §A.3).
Data Sampler. Training retrieval models typically require both positive and negative data pairs. The positive pair naturally consist of the NL description and pseudocode function from the same source code. To construct effective negative samples, we employ a hybrid strategy combining random sampling and hard negative mining hardnegativemining2025CIKM. Specifically, for each source code function, we take the NL description as a query and randomly select a bunch of binary functions as negative candidates that are compiled from the other source code among the whole corpus. Furthermore, we utilize a general embedding model (Qwen3-Embedding-8B 2025qwen3-emb), which is top-tier in the MTEB leaderboard 2025mmteb, to estimate the semantic similarity between the query and the NL descriptions of the negative candidates. After arranging in descending order of semantic similarity, we select the first binary function with a score below the threshold ($\leq 0.95$) as a hard negative sample. In this way, we constructed 45,703,009 high-quality positive and negative data pairs in total, enabling the model to learn the mappings across different semantic representations.
@@ 3.2 Model Architecture
BinSeek-Embedding. The embedding model is built from scratch with reference to the existing language model Qwen3 qwen3. Specifically, we initialize our model backbone with an 8 transformer layer, which incorporates RMSNorm RMSNorm, SwiGLU activation SwiGLU, and grouped query attention (GQA) GQA techniques for improved training stability and performance. We employ Qwen3’s tokenizer that implements byte-level byte-pair encoding Wang_Cho_Gu_2020; sennrich-etal-2016-neural with a vocabulary size of 151,669. Consequently, the weights of the token embedding layer are directly loaded from the well-trained Qwen3 model and frozen to reduce training overhead. Following the embedding layer, we use a rotary position embedding (RoPE) RoPE layer to provide position information to transformer blocks, meanwhile enabling input length scaling at inference time. Finally, the last token pooling is employed to the backbone output to generate the final embedding.
@@
To mitigate this, we here propose a heuristic function importance metric to select the top-$k$ important callees to provide context. Existing studies have demonstrated that the presence of meaningful function names and strings often provides significant semantic cues about its functionality NER-2023pst; BinMetric; BinSum-jin2023. Therefore, we try to measure the meaningful symbol density of a pseudocode function $f_{t}$ through three aspects: (1) whether the function has an unstripped function name, (2) the density of strings among the pseudocode, and (3) the meaningful callee names in the function. Formally, we define the informative score $\mathcal{S}(f_{t})$ to indicate function importance as follows: